# Evaluating and Steering Modality Preferences in Multimodal Large Language Model

### v2

@@ 4.1 Which Modality Do MLLMs Prioritize?
We use the MC2 benchmark to evaluate the modality preferences of 20 open-source MLLMs and the proprietary ChatGPT-4o-mini (Hurst et al., 2024a), detailed in Appendix C.1.
Different MLLMs exhibit different modality preferences. As described in Section 3.3, we quantify modality preference using Vision Ratio, with the results presented in the left panel of Figure 2 and detailed in Table 6. We observe that all MLLMs exhibit clear modality preference, with most models showing a strong preference for text; for instance, LLaVA1.5-7B attains only a 13.4% Vision Ratio. This aligns with the previous findings that MLLMs suffer from a severe language prior (Lee et al., 2024; Parcalabescu & Frank, 2024; Wu et al., 2025). Interestingly, the Qwen2.5VL and InternVL3 show a certain degree of preference towards the vision modality.
Larger MLLMs exhibit stronger preferences for the vision modality. We evaluate models from the LLaVA1.5, LLaVA-Next, Qwen2.5VL, InternVL3, and LLaVA-OneVision families to investigate the relationship between model size and modality preference. As shown in the right panel of Figure 2, we observe that for all model families, the preference for the vision modality increases with the model size. And the Qwen2.5VL and InternVL3 models exhibit a significant preference for the vision modality once the model size increases. However, LLaVA1.5, LLaVA-Next, and LLaVA-OneVision models maintain a noticeable preference for the text modality as their sizes increase.
Vision Ratio aligns with human preference. We further verify whether the Vision Ratio can serve as a human-level measure of modality preference in MLLMs. We randomly sample 100 instances from MC2 and compute the Vision Ratio of four representative models—Qwen2.5VL-7B, LLaVA-OneVision-7B, InternVL3-14B, and LLaVA1.5-7B. In addition, we craft prompts to elicit explicit reasoning chains from the models, specifically targeting their reliance on visual or textual information. To ensure labeling reliability, three expert annotators independently annotate the expressed modality preference for each response, with the final label determined by majority vote. The automatically obtained Vision Ratio scores ([56.3%, 24.6%, 52.3%, 13.9%]) are highly consistent with the ones given by human ([61.0%, 22.0%, 51.0%, 16.0%]), with an average discrepancy of only 2.68%. This indicates that the Vision Ratio can act as a reliable, automated proxy for human assessment of modality preference.
Figure 3: Analysis of modality preferences. Left: Trends of Vision Ratio and multimodal Attention Ratio across different models. Middle: Vision Ratio with Respect to the Proportion of Multimodal Conflict-Context Training Data and different MLLMs. Right: Relationship between visual understanding ability, quantified as the average accuracy across seven widely used benchmarks and the modality preference measured by Vision Ratio.
@@ 4.2 What Factors Drive These Preferences?
Given that different MLLMs display varying modality preferences, we examine two primary sources of such differences: the internal attention distribution and the training factors.
Different allocation of attention across modalities. We compute the mean attention scores over all token positions from both modalities and define the ratio of visual attention to total attention as the Attention Ratio. By analyzing Qwen2.5VL-7B and LLaVA-OneVision-7B, we observe that the trends of the Attention Ratio closely align with the Vision Ratio across models in the left panel of Figure 3. This alignment suggests that MLLMs distribute attention unevenly between modalities, which in turn contributes to their divergent modality preferences.
Impact of model scale and training data recipe. Through reviewing the technical reports of the evaluated MLLMs, we find that they all adopt a common architecture comprising a vision encoder, an alignment layer, and an LLM. Thus, we hypothesize that the observed preferences mainly arise from two factors: 1) Exposure to more multimodal contexts, especially with conflicting cases, drives more pronounced shifts in modality preference. 2) Larger LLMs are more capable of shifting their preference during training;
To examine these hypotheses, we construct a training dataset containing vision–text conflict contexts and fine-tune Qwen2.5VL-7B/3B and LLaVA1.5-7B with varying proportions of samples with multimodal conflict contexts, adjusting their preferences toward text or vision. We then measured changes with the Vision Ratio. We optimize MLLMs in the opposite direction of their original preferences and measure changes using the Vision Ratio. As shown in the middle panel of Figure 3, increasing the proportion of multimodal contexts consistently leads to larger preference shifts, supporting Hypothesis 1. This suggests that multimodal inputs maybe create more challenging training conditions, leading to stronger shifts of preference. Furthermore, Qwen2.5VL-7B exhibites greater shifts than Qwen2.5VL-3B under the same conditions, supporting Hypothesis 2. This indicates that larger LLMs demonstrate stronger learning ability and adapt more effectively.
@@ 4.3 Can the Vision Ratio provide guidance for downstream task performance?
To demonstrate the correlation between the modality preference and performance for downstream tasks, we evaluate the visual understanding abilities of 10 representative MLLMs. Specifically, we compute the average accuracy across 7 widely used benchmarks, as detailed in the Appendix C.2. We then compare the visual understanding abilities with their modality preference measured by Vision Ratio using MC2, as shown in the right panel of Figure 3. The results reveal a strong positive association between the Vision Ratio and visual understanding ability across the evaluated MLLMs. Specifically, Spearman’s rank correlation (Sedgwick, 2014) reaches $\rho=0.964$, demonstrating that the Vision Ratio provides a highly reliable indicator of visual understanding task performance.
@@ 4.4 Can Modality Preference Be Controlled?
We employ instruction guidance to investigate whether modality preference can be controlled, and conduct a latent space representation analysis to examine the mechanisms, underlying the preference adjustment. Details of the experimental design and results are provided in Appendix C.3.
Figure 4: Analysis of modality preference under instruction-guidance. Left: Adjustment of modality preference using instruction-guided control (Inst-vision vs. Inst-text). Middle: Representation shifts under instruction-guided interventions. Right: layer-wise absolute difference and standard deviation of hidden states between different instruction.
Modality preference can be guided through instruction design. We investigate the impact of instruction design on modality preference, for instance, by explicitly directing the model to rely on a specific modality when answering a question. Specifically, we evaluate the modality preference measured by Vision Ratio for four representative MLLMs including Qwen2.5VL-7B, LLaVA-OneVision-7B, Qwen2VL-7B and InternVL3-14B under the text or vision preference instruction (Inst-text, Inst-vision). As illustrated in the left panel of Figure 4, instructions that steer the model toward a particular modality effectively shape its modality preference.
Modality preference direction in representation space. To further understand how the intervention methods influence modality preference internally, we analyze the hidden representations of the models. Specifically, we apply Principal Component Analysis (PCA; Abdi & Williams, 2010) to the hidden states to identify the dominant direction corresponding to modality preference shifts. The middle panel of Figure 4 shows that instruction-based interventions drive clear shifts in representations, aligning with the modality specified by the instruction. The PCA direction further reveals that the model’s internal states are sensitive to modality control cues, which motivates us to develop representation techniques for adjusting modality preference and to apply these insights to downstream tasks in Section 5.
@@
5 Representation Based Modality Preference Steering
Inspired by the representation behavior discussed in Section 4.4, we propose to use the representation engineering (Zou et al., 2023) to steer the modality preference, controlling the model’s behavioral expression. As shown in Figure 5, the proposed framework consists of Modality Preference Probing (§5.1) and Modality Preference Steering (§5.2).
@@ 5.1 Modality Preference Probing
We probe and collect neural activity that represents the direction of modality preference. Inspired by the pre-training Next Token Prediction objective of decoder-only MLLMs (Hurst et al., 2024a) and the method to extract classification features (Feucht et al., 2024), we collect neural activity from the last token in the input text. The process involves probing modality preference through two pairs of requests: one with a vision preference probing (e.g., ‘answer the question based on the vision context’) and another with a text preference probing (e.g., ‘answer the question based on the text context’). Let us denote these two inputs by $q^{v}$ (based on the vision context) and $q^{t}$ (based on the text context), and consider a set of $N$ such pairs ($q^{v}_{i},q^{t}_{i}$), $i\in\{1,\dots,N\}$. Let $\mathbf{x}_{i,\ell}^{v},\mathbf{x}_{i,\ell}^{t}\in\mathbb{R}^{d}$ be the hidden states on the two queries at the last token of the input at layer $\ell\in\{1,\dots,L\}$, where $d$ is the dimension of the chosen MLLM. We identity the direction of modality preference by computing the difference in the hidden states between the paired inputs. More formally, we compute a vector $\mathbf{u}_{\ell}\in\mathbb{R}^{d}$ representing the direction towards the text modality at layer $l$ for a given query as:
@@
Averaging over different queries allows us to capture the activation values most closely associated with modality preference, independent of questions. As shown in the right panel of Figure 4, we compute the absolute values and standard deviations of the modality preference direction $\mathbf{u}^{t}_{\ell}$ across different samples. We observe that layers 20–23 exhibit both higher absolute values and lower variance, indicating that the preference direction is more prominent and stable in these layers. Based on this observation, we select the corresponding layer $\ell^{\prime}$ of the model to control the direction of modality preference in Section 5.2. Similar patterns are observed for Qwen2VL-7B, Qwen2.5VL-7B, LLaVA-OneVision and InternVL3, as detailed in Appendix D.1.
5.2 Modality Preference Steering
After obtaining the probing direction vector, we compute the steering vector by re-scaling the vector $\mathbf{u}^{t}_{\ell}$ with a weight $w\in\mathbb{R}_{d}$. The scaling process must carefully balance two objectives: 1) it must be strong enough to effectively steer the model’s modality preference, 2) it must preserve the model’s normal output behavior. In our preliminary experiments, we observe that setting the weight too large leads to repetitive and meaningless outputs, whereas a too small weight fails to obtain any noticeable change for modality preference. Unlike previous approaches (Zou et al., 2023; Stolfo et al., 2024) that rely on exhaustive search over a validation set to determine the weight, we propose a principled method that aligns the mean of the probed direction distribution with the mean of the original distribution of hidden states. This strategy ensures that the steering remains effective without disrupting the model’s inherent generation capabilities. Formally, the weight is determined by aligning the mean of the probed direction with the central distribution of the original hidden states and the steering vector is computed by:
Finally, during the second inference, we adjust the original hidden states at the selected steering layer $\ell^{\prime}$ at each decoding step by adding $\mathbf{s}^{t}_{\ell^{\prime}}$ to all tokens to steer the model’s response toward the text modality. Similarly, steering models towards vision modality is performed towards the opposite direction. In the final implementation, no additional data or labels are introduced. We only require two consecutive rounds of inference: the first for probing and the second for steering, effectively controlling the modality preference.
Table 1: Performance for steering Qwen2VL-7B and OneVision-7B towards vision modality and steering Qwen2.5VL-7B and InternVL3-8B towards text modality, measured by $S_{vision}$ and $S_{text}$.
@@ 5.2 Modality Preference Steering
InternVL3-8B 37.5 52.8 53.1 49.8 57.1
Table 2: Multimodal translation results for Ambigcaps (Li et al., 2021). BLEU scores are reported for English (En) $\leftrightarrow$ Turkish (Tr).
Method En->Tr Tr->En
Qwen2.5VL-7B 8.92 18.56
      +Inst towards vision  8.21 (-0.71) 16.09 (-2.47)
      +Inst towards text  9.45 (+0.53) 18.98 (+0.42)
      +Ours 10.22 (+1.30) 19.89 (+1.33)
@@ 5.3 Experiments
We verify the effectiveness of the method in controlling modality preference on MC2 and downstream tasks across Qwen2VL-7B, Qwen2.5VL-7B, and LLaVA-OneVision-7B and InternVL3-8B. We consider widely used training-free approaches as baselines: MLLM refers to employing MLLM to directly reason in modality-conflicting contexts; InstDesign uses instructions to guide modality preference direction; CoT enables complex reasoning through intermediate steps; and FewShot uses four examples to guide the models. For detailed implementation and results, refer to Appendix D.
As shown in Table 2, the proposed method consistently outperforms the baseline approaches on MC2 across both settings, demonstrating its effectiveness in adjusting modality preference. We further assess the effectiveness of the proposed method on two types of downstream tasks: 1) multimodal machine translation (MMT) using AmbigCaps (Li et al., 2021), and 2) visual understanding on PhD (Liu et al., 2024c). The latter includes two subsets—PhD-ica, which contains irrelevant textual context, and PhD-icc, which introduces misleading or incorrect textual information—both of which increase the risk of hallucination. In MMT, the task should primarily ground the source-language text while treating visual information as auxiliary. Accordingly, we adjust modality preference toward the text modality. As shown in Table 2, our method yields an improvement of 1.33 BLEU score over the baselines. By contrast, for PhD, we steer the modality preference toward the vision modality to ground the image. The results in Table 3 demonstrate that our approach achieves substantial improvements across both MLLMs. In particular, when applied to Qwen2VL-7B, our method surpasses the best baseline by an average margin of 6.1 percentage points.

### v3

@@ 4.1 Which Modality Do MLLMs Prioritize?
We use MC2 to evaluate the modality preferences of 20 open-source MLLMs and the proprietary ChatGPT-4o-mini (Hurst et al., 2024a), detailed in Appendix C.1.
Different MLLMs exhibit different modality preferences. As described in Section 3.3, we quantify modality preference using Vision Ratio in the left panel of Figure 2 and detailed in Table 13. We observe that all MLLMs exhibit clear modality preference, with most models showing a strong preference for text; for instance, LLaVA1.5-7B attains only a 13.4% Vision Ratio. This aligns with the previous findings that MLLMs suffer from a severe language prior (Lee et al., 2024; Parcalabescu and Frank, 2024; Wu et al., 2025). Interestingly, Qwen2.5VL and InternVL3 show a certain degree of visual preference.
Larger MLLMs exhibit stronger visual preferences. We evaluate models from the LLaVA1.5, LLaVA-Next, Qwen2.5VL, InternVL3, and LLaVA-OneVision families to investigate the relationship between model size and modality preference. As shown in the right panel of Figure 2, we observe that for all model families, the preference for the vision modality increases with the model size. And Qwen2.5VL and InternVL3 models exhibit a significant preference for the vision modality once the model size increases. However, other models maintain a noticeable textual preference as model sizes increase. To validate the reliability of the evaluation, we conduct a sensitivity analysis for the sample number of MC2 in Appendix E.1.
Vision Ratio aligns with human preference. We further verify whether the Vision Ratio can serve as a human-level measure of modality preference in MLLMs. We randomly sample 100 instances from MC2 and compute the Vision Ratio of four representative models—Qwen2.5VL-7B, OneVision-7B, InternVL3-14B, and LLaVA1.5-7B. In addition, we craft prompts to elicit explicit reasoning chains from the models, specifically targeting their reliance on visual or textual information. To ensure labeling reliability, three expert annotators independently annotate the expressed modality preference for each response, with the final label determined by majority vote. The automatically obtained Vision Ratio scores ([56.3%, 24.6%, 52.3%, 13.9%]) are highly consistent with those given by human annotators ([61.0%, 22.0%, 51.0%, 16.0%]), with an average discrepancy of only 2.68%. This indicates that the Vision Ratio can act as a reliable, automated proxy for human assessment of modality preference.
@@ 4.2 What Factors Drive These Preferences?
Given that different MLLMs display varying modality preferences, we examine two primary sources: the internal attention distribution and the training factors.
Different allocation of attention across modalities. We compute the mean attention scores over all token positions from both modalities and define the ratio of visual attention to total attention as the Attention Ratio. By analyzing Qwen2.5VL-7B and OneVision-7B, we observe that the trends of the Attention Ratio closely align with the Vision Ratio across models in the left panel of Figure 3. This alignment suggests that MLLMs distribute attention unevenly between modalities, which in turn contributes to their divergent modality preferences.
Figure 4: Analysis of modality preference under instruction-guidance. Left: Adjustment of modality preference using instruction-guided control (Inst-vision vs. Inst-text). Middle: Representation shifts under instruction-guided interventions. Right: Layer-wise absolute difference and standard deviation of hidden states between different instructions.
Impact of model scale and training data recipe. Through reviewing the technical reports of the evaluated MLLMs, we find that they all adopt a common architecture comprising a vision encoder, an alignment layer, and an LLM. Thus, we hypothesize that the observed preferences mainly arise from two factors: 1) Exposure to more multi-modal contexts, especially with conflicting cases, drives more pronounced shifts in modality preference; 2) Larger LLMs are more capable of shifting their preference during training.
To examine these hypotheses, we construct a training dataset containing vision–text conflict contexts and fine-tune Qwen2.5VL-7B/3B and LLaVA1.5-7B with varying proportions of samples with multi-modal conflict contexts to adjust preferences toward text or vision. We optimize MLLMs in the opposite direction of their original preferences and measure changes using the Vision Ratio. As shown in the middle panel of Figure 3, increasing the proportion of multi-modal contexts consistently leads to larger preference shifts, supporting Hypothesis 1. This suggests that multi-modal inputs may create more challenging conditions, leading to greater preference shifts. Furthermore, Qwen2.5VL-7B exhibits greater shifts than Qwen2.5VL-3B under the same training data, supporting Hypothesis 2. This indicates that larger LLMs demonstrate stronger adaptability.
@@ 4.3 Can the Vision Ratio provide guidance for downstream task performance?
As a foundational behavioral prior, the identified modality preference can inform how a model integrates information across modalities. As such, the findings can offer relevant insights into the model’s behavior in deeper cross-modal understanding tasks. To demonstrate the correlation between the modality preference and performance on downstream tasks, we evaluate the visual understanding abilities of 10 representative MLLMs. Specifically, we compute the average accuracy across 7 widely used benchmarks, as detailed in Appendix C.2. We then compare the visual understanding abilities with their modality preference measured by Vision Ratio using MC2, as shown in the right panel of Figure 3. The results reveal a strong positive association between Vision Ratio and visual understanding ability across all MLLMs. Notably, Spearman’s rank correlation (Sedgwick, 2014) reaches $\rho=0.964$, demonstrating that Vision Ratio provides a highly reliable indicator of visual understanding ability.
@@ 4.4 Can Modality Preference Be Controlled?
We investigate whether modality preference can be controlled, and conduct latent space representation analysis to examine the underlying mechanisms. The detailed setting and results are provided in Appendix C.3.
Modality preference can be guided through instruction design. We investigate the impact of instruction design on modality preference, for instance, by explicitly directing the model to rely on a specific modality when answering a question. Specifically, we evaluate the modality preference using Vision Ratio for four representative MLLMs including Qwen2.5VL-7B, OneVision-7B, Qwen2VL-7B and InternVL3-14B under the text and vision preference instruction (Inst-text and Inst-vision). As shown in the left panel of Figure 4, instructions, probing the models toward any modality effectively shape their modality preferences.
Modality preference direction in representation space. To further understand how the intervention methods influence modality preference internally, we analyze the hidden representations of the models. Specifically, we apply Principal Component Analysis (PCA; Abdi and Williams, 2010) to the hidden states to identify the dominant direction of modality preference. The middle panel of Figure 4 shows that instruction-based interventions drive clear shifts in representations, revealing that the model’s internal states are sensitive to modality control. This motivates us to develop specific techniques to adjust modality preference in Section 5.
@@
5 Method
Inspired by the behavior discussed in Section 4.4, we propose to use representation engineering (Zou et al., 2023) to steer modality preference. As shown in Figure 5, the proposed framework consists of Modality Preference Probing (§5.1) and Modality Preference Steering (§6).
@@ 5.1 Modality Preference Probing
We probe and collect neural activity that represents the direction of modality preference. Inspired by the pre-training Next Token Prediction objective of decoder-only MLLMs (Hurst et al., 2024a) and the method to extract classification features (Feucht et al., 2024), we collect neural activity from the last token in the input text. The process involves probing modality preference through two requests: one with a vision preference probing (e.g., ‘answer the question based on the vision context’) and another with a text preference probing (e.g., ‘answer the question based on the text context’). Let us denote these inputs by $q^{v}$ and $q^{t}$, and consider a set of $N$ such pairs ($q^{v}_{i},q^{t}_{i}$), $i\in\{1,\dots,N\}$. Let $\mathbf{x}_{i,\ell}^{v},\mathbf{x}_{i,\ell}^{t}\in\mathbb{R}^{d}$ be the hidden states for the two queries at the last token at layer $\ell\in\{1,\dots,L\}$, where $d$ is the dimension of the chosen MLLM. We identify the direction of modality preference by computing the difference in the hidden states between the paired inputs. Formally, we compute a vector $\mathbf{u}_{\ell}\in\mathbb{R}^{d}$ representing the direction towards the text modality at layer $\ell$ for a given query as:
@@
Table 1: Performance for steering Qwen2VL-7B and OneVision-7B towards vision modality, and Qwen2.5VL-7B and InternVL3-8B towards text modality, measured by $S_{\text{vision}}$ and $S_{\text{text}}$.
@@
OneVision-7B 37.5 52.8 53.1 49.8 57.1
Averaging over different queries allows us to capture the activation values most closely associated with modality preference, independent of questions. As shown in the right panel of Figure 4, we compute the absolute values and standard deviations of the modality preference direction $\mathbf{u}^{t}_{\ell}$ across different samples. We observe that layers 20–23 exhibit higher absolute values and lower variance, indicating that the preference direction is more prominent and stable in these layers. Based on this observation, we select the corresponding layer $\ell^{\prime}$ of the model to control the direction of modality preference in Section 6. Similar patterns are observed for Qwen2VL-7B, Qwen2.5VL-7B, LLaVA-OneVision and InternVL3, as detailed in Appendix D.1.
Table 2: Performance of the proposed method on the reasoning tasks (MathVista, TallyQA and VSR) measured by accuracy for Qwen2.5VL-7B and multi-modal translation task (Ambigcaps) measured by BLEU scores for English (En) $\leftrightarrow$ Turkish (Tr).
Dataset CoT InstDesign Ours
MathVista 50.0 59.0 62.3
TallyQA 61.6 74.4 76.6
VSR 45.6 51.3 53.2
Ambigcaps (En→Tr) 9.03 9.45 10.2
Ambigcaps (Tr→En) 18.63 18.98 19.89
@@
5.2 Modality Preference Steering
Figure 6: The attention scores of the generated token toward the vision and text contexts at the 22th layer using InstDesign (Left) and the proposed method (Right).
After obtaining the probing direction vector, we compute the steering vector by re-scaling the vector $\mathbf{u}^{t}_{\ell}$ with a weight $w$. The scaling process must carefully balance two objectives: 1) it must be strong enough to effectively steer the model’s modality preference, 2) it must preserve the model’s normal output behavior. In our preliminary experiments, we observe that setting the weight too large leads to repetitive and meaningless outputs, whereas a too small weight fails to obtain any noticeable change for modality preference. Unlike previous approaches (Zou et al., 2023; Stolfo et al., 2024) that rely on exhaustive search over a validation set to determine the weight, we propose a principled method that aligns the mean of the probed direction distribution with the mean of the original distribution of hidden states. This strategy ensures that the steering remains effective without disrupting the model’s inherent generation capabilities. Formally, the weight is determined by aligning the mean of the probed direction with the central distribution of the original hidden states and the steering vector is computed by:
Finally, during the second inference, we adjust the hidden states at the selected layer $\ell^{\prime}$ at each decoding step by adding $\mathbf{s}^{t}_{\ell^{\prime}}$ to all tokens to steer the model’s response toward text modality. Similarly, steering towards vision modality is performed towards the opposite direction and no additional data or labels are introduced.
@@ 5.3 Experiments
We evaluate the proposed method for controlling modality preference on MC2 and multiple downstream tasks across four representative MLLMs: Qwen2VL-7B, Qwen2.5VL-7B, OneVision-7B, and InternVL3-8B. We compare our approach against several widely-used training-free baselines: MLLM (zero-shot direct response), InstDesign (instruction-based guidance for modality preference), CoT (utilizing intermediate processing steps), and FewShot (four-shot in-context learning). Implementation details are provided in Appendix D. As shown in Table 1, our method consistently outperforms all baselines on MC2 across both evaluation settings. These results demonstrate the superior effectiveness of our approach in precisely adjusting modality preference compared to standard prompting techniques.
We further assess our method across a variety of downstream scenarios: 1) Multi-modal reasoning tasks on MathVista (Lu et al., 2023), TallyQA (Acharya et al., 2019) and VSR (Liu et al., 2023a); 2) Multi-modal understanding tasks on PhD (Liu et al., 2024c) and AmbigCaps (Li et al., 2021), a multi-modal machine translation (MMT) task. The Phd dataset includes two subsets—PhD-ica, which contains irrelevant textual context, and PhD-icc, which introduces misleading or incorrect textual information—both of which increase the risk of hallucination.
As shown in Table 2, our method consistently surpasses both the CoT and InstDesign baselines. We observe that by steering the model toward a visual preference, our method increases reliance on the original visual input. This effectively prevents the final decision from being misled by potential vision hallucination that often emerges during the sequential steps of CoT. On the AmbigCaps translation task, our method achieves a significant improvement of $1.33$ BLEU score over the strongest baseline. Furthermore, Table 3 demonstrates substantial improvements in multi-modal understanding. Specifically, when applied to Qwen2VL-7B, our method surpasses the best-performing baseline by an average margin of $6.1$ percentage points. Additional results for more tasks and models are available in Appendix E.2.
5.4 In-depth Analysis for Steering Method
To investigate the internal mechanism of the steering method, we analyze the attention scores of the generated token toward the vision and text contexts using the proposed steering method and the InstDesign method with the samples from $MC^{2}$. In Figure 6, we visualize the attention distribution at the 24th layer by steering the Qwen2.5VL-7B at the 22nd layer towards the text modality using the case in Figure 12, as detailed in Appendix E.4. We observe that after applying the steering method, the model’s attention weight toward the text modality significantly increases. This change clearly demonstrates that the steering mechanism successfully alters the modality preference by enhancing the model’s dependency on the text modality. More ablation experiments, the analysis of the latency, memory usage and the prerequisites for the proposed method are detailed in Appendix E.3 and  E.4.